OpenStudy (anonymous):

Find the directrix of the parabola with the equation y+3=(1/10)[(x+2)^2]

OpenStudy (amistre64):

well, 4p = 10 will help

OpenStudy (anonymous):

?

OpenStudy (amistre64):

the standard geometric equation for a parabola such as y=x^2, is of the form: 4py = x^2 ; there p is a distance that measures the difference from the vertex to the directrix, or the vertex to the focus

OpenStudy (amistre64):

y = 1/10 x^2 10y = x^2 4p = 10 will help us define the distance we need to move from the vertex

OpenStudy (amistre64):